OpenStudy (anonymous):

How do I find the zeroes for this problem? (x+5)^2/(x-3)^2+(x+5)^3/(x-3)^3 = 0

OpenStudy (precal):

x=-5 I believe you are solving just the numerators

OpenStudy (anonymous):

you have to combine the fractions and make it one. then find the x values that will make the numerator part 0.

OpenStudy (anonymous):

-1 too, I think

OpenStudy (precal):

I did a quick graph and it shows -5 as a zero and -1

OpenStudy (mathmate):

(x+5)^2/(x-3)^2+(x+5)^3/(x-3)^3 = 0 [ (x+5)^2(x-3) + (x+5)^3 ] /(x-3)^3 = 0 Factor the numerator to get (x+5)^2(2x+2)=0 => x=-5 (multiplicity 2) and x=-1 as precal got.

OpenStudy (anonymous):

Thanks! I was so very close...